# Gold after Randomized Sand: Model-X Split Knockoffs for Controlled Transformation Selection

### v1

@@ 4 Simulation Experiments
In this section, the performance of the proposed Model-X Split Knockoff method is evaluates in two canonical settings, the normal distribution setting and the pairwise comparisons setting. The simulation results validates that the Split Knockoff method achieves desired false discovery rate control as well as relatively high selection power, specifically higher compared with Model-X Knockoff when both are applicable.
@@ 4.1 Normal Distribuions
Experimental Setting: In this section, simulation experiments are conducted under the case where the distribution of the design $X$ follows the normal distribution. In particular, rows of the design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ is generated independent and identically distributed (i.i.d.) from a multivariate normal distribution $\mathcal{N}(0_{p},{\bm{\Sigma}})$, where the covariance matrix ${\bm{\Sigma}}$ has unit diagonal entries (${\bm{\Sigma}}_{i,i}=1$) and correlation structure ${\bm{\Sigma}}_{i,j}=c^{|i-j|}$ for $i\neq j$ with correlation parameter $c=0.5$:
${\bm{\Sigma}}=\begin{pmatrix}1&c&c^{2}&c^{3}&\cdots&c^{p-1}\\ c&1&c&c^{2}&\cdots&c^{p-2}\\ c^{2}&c&1&c&\cdots&c^{p-3}\\ c^{3}&c^{2}&c&1&\cdots&c^{p-4}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ c^{p-1}&c^{p-2}&c^{p-3}&c^{p-4}&\cdots&1\end{pmatrix}.$
The true coefficient vector ${\beta}^{*}\in\mathbb{R}^{p}$ is defined by
$\beta^{*}_{i}=\begin{cases}A&i\leq k,\quad i\equiv 0,-1\pmod{3},\\ 0&\text{otherwise}.\end{cases}$ (39)
In this section, we consider the logistic regression, where the response $\mathbf{Y}\in\mathbb{R}^{n}$ is generated from
$\mathbf{Y}\sim\text{Bernoulli}\left\{\sigma(\mathbf{X}\beta^{*})\right\}.$ (40)
Here, $\sigma(x)$ denotes the sigmoid function, i.e. $\sigma(x)=\frac{1}{1+\exp(-x)}$.
@@
The transformational sparsity is specified by the linear transformation matrix $D\in\mathbb{R}^{m\times p}$ such that $\gamma^{*}=D\beta^{*}$ is sparse. Based on the above specific choice of $\beta^{*}$, the following three kinds of $D$ is considered in this simulation:
  • $\beta^{*}$ itself is sparse, so that one can take $D_{1}=I_{p}$, where $I_{p}$ is the identity matrix of size $p\times p$, leading to $m=p$.
@@
  • $\beta^{*}$ is a uni-dimensional piecewise constant function, so that one can take $D_{2}$ as the 1-D graph difference operator on a line. Specifically, $D_{2}\in\mathbb{R}^{(p-1)\times p}$, where $D_{2}(i,i)=1$, $D_{2}(i,i+1)=-1$ for $i\in\{1,\dots,p-1\}$, and $D_{2}(i,j)=0$ for other pairs of $(i,j)$. In this case, $m=p-1<p$.
@@
  • Combining the above two cases, one can take $D_{3}=\begin{pmatrix}D_{1}\\ D_{2}\end{pmatrix}\in\mathbb{R}^{(2p-1)\times p}$, where $m=2p-1>p$.
@@
For the case where $D_{1}=I_{p}$, the performance of Model-X Knockoffs is also evaluated for comparisons. Throughout this section, the target FDR level $q$ is set to be $0.2$, while the coefficients in the above simulation design are taken to be $p=100$, $k=20$, and $A=1$. The Model-X Knockoff method is implemented using the official repository [^4] https://github.com/msesia/knockoff-filter.
The Model-X (Split) Knockoff copies are generated from the data matrix $\mathbf{X}$ through the estimated covariance matrix $\widehat{\bm{\Sigma}}=\mathbf{X}^{T}\mathbf{X}/n$. Specifically, the Split Knockoff copies $A$ and $\widetilde{A}$ are generated through Equation (25) and (20). To select the tuning parameters $\lambda$ and $\nu$ in Equation (2.3), cross-validation is applied to identify the optimal values, $\widehat{\lambda}$ and $\widehat{\nu}$, based on the test loss. To solve the optimization problem (2.3), we leverage the Embedded Conic Solver (ECOS) algorithm for convex optimization with the CVX [^5] https://www.cvxpy.org package.
4.1.1 Comparisons with Model-X Knockoffs when $D=I_{p}$
(i) Performance of Model-X Knockoff
(ii) Performance of Model-X Split Knockoff
Figure 1: Performance of Model-X Knockoffs and Model-X Split Knockoffs when the number of samples varies in the case $D=D_{1}=I_{p}$. The logistic regression is implemented for $q=0.2$. The curves in the figures represent average performance and the shaded areas represent the 80% confidence intervals in 200 simulations.
In Figure 1, we compare the performance of Model-X Knockoffs and Model-X Split Knockoffs when $D=D_{1}=I_{p}$, i.e. the case where the transformation sparsity problem is degenerated into the classical direct sparsity problem. In both figures, with the increase of sample size, the selection power of both methods gradually increases from 0 to 1, while the FDR of both methods are under control for all sample sizes.
@@ 4.1.1 Comparisons with Model-X Knockoffs when $D=I_{p}$
In terms of selection power, the Model-X Split Knockoff demonstrably yields superior results across all tested sample sizes, a finding supported by Proposition 2.1. We hypothesize that this performance gain is a direct consequence of the improved incoherence conditions, which arise from the orthogonal design of $A$ and $\widetilde{A}$ in Equations (25) and (30).
4.1.2 Performance of Model-X Split Knockoffs on Non-trivial $D$
(i) $D=D_{2}$
(ii) $D=D_{3}$
Figure 2: Performance of Model-X Split Knockoff when the number of samples varies in the cases where $D=D_{2}$ or $D=D_{3}$. The logistic regression is implemented for $q=0.2$. The curves in the figures represent average performance and the shaded areas represent the 80% confidence intervals in 200 simulations.
@@ 4.1.2 Performance of Model-X Split Knockoffs on Non-trivial $D$
In Figure 2, the performance of Model-X Split Knockoffs is evaluated in the scenarios where $D$ is non-trivial, i.e. $D\neq I_{p}$. In both cases, the FDR of Model-X Split Knockoffs is under control universally for all choices of sample sizes. Meanwhile, the selection power of Model-X Split Knockoffs follows a similar trend as it is in the case where $D=I_{p}$, i.e. the selection power gradually increases from 0 to 1 when the sample size enlarges.
@@ 4.2 Pairwise Comparisons
Experimental Settings: In this section, the pairwise comparison problem among $p$ objects with $n$ noisy observations of comparisons is considered. The scores for $p$ objects, denoted as $\beta^{*}\in\mathbb{R}^{p}$, are generated from a standard normal distribution $\mathcal{N}(0,1)$. To introduce sparsity into the dataset, the scores in $\beta^{*}$ are truncated according to varying sparsity ratios $k$ set at 0.1, 0.3, and 0.5. Specifically, only the first $k\cdot p$ components of $\beta^{*}$ are retained, while the remaining components are set to zero.
The rows of the pairwise comparison design matrix $\mathbf{X}\in\{0,1,-1\}^{n\times p}$ is generated independently in the following way. For the $i$-th row, two indices $1\leq j<k\leq p$ are chosen uniformly at random with probability $\frac{2}{p(p-1)}$ for each possible pair of $(j,k)$. Then we set $\mathbf{X}_{i,j}=1$, $\mathbf{X}_{i,k}=-1$, and $\mathbf{X}_{i,l}=0$ for $l\neq j,k$. This specific construction ensures that each row of $\mathbf{X}$ represents a (noisy) comparison between two objects.
In this simulation experiment, the performance of Model-X Split Knockoff is evaluated under the Bradley-Terry model, where the response vector $\mathbf{Y}\in\mathbb{R}^{n}$ is generated by Equation (40). The transformational matrix $D\in\mathbb{R}^{\frac{p(p-1)}{2}\times p}$ is taken to be the graph difference operator on a full connected graph with $p$ vertices to conduct pairwise comparisons.
@@
Figure 3: Performance of Model-X Split Knockoff as the Number of Features Varies for Bootstrap+ and Sequential Constructions. Simulation experiments are conducted using the Bradley-Terry Model of pairwise comparison with $q=0.2$ and a sparsity level of $0.5$ for $\beta^{*}$ (i.e., $k/p$ in Equation (39)). The curves represent average performance, while the shaded areas indicate 80% confidence intervals based on 200 simulation runs.
Figure 3 compares the performance of bootstrap+ constructions of Model-X Split Knockoff copies against the sequential constructions. It shows that, although the two types of constructions achieve desired FDR control, the selection power of the bootstrap+ constructions is higher compared with the sequential constructions when the number of features varies. This is explained in Section 2.2.2 that the bootstrap+ constructions forces $\widetilde{X}^{A}$ to be independent from $A$, which helps improve the incoherence conditions and thus the selection power.
Figure 4: Runtime(Second) of Model-X Split Knockoff for bootstrap+ and sequential methods. Results are averaged on 200 simulation runs with the number of samples $n=1000$, and the number of features varying from 5 to 30.
Moreover, when $p$ is large, the sequential construction may need to estimate too many conditional distributions, that the computation cost explodes. On the other hand, the bootstrap+ constructions does not suffer from this issue, and handles such a case easily. As shown in Figure 4, the sequential method exhibits a much steeper increase in running time compared to the bootstrap+ method, which requires significantly less computation time under the same settings.
Due to the above two points, the bootstrap+ construction will be the default construction to implement Model-X Split Knockoff for pairwise comparisons in the following sections.
(i) Sparsity Level $k=0.1$
(ii) Sparsity Level $k=0.3$
(iii) Sparsity Level $k=0.5$
Figure 5: Performance of Model-X Split Knockoff when the number of samples varies. Simulations experiments are implemented through Bradley-Terry Model of pairwise comparison for $q=0.2$ with different sparsity level of $\beta^{*}$. The curves in the figures represent average performance and the shaded areas represent the 80% confidence intervals in 200 simulations.
Figure 5 illustrate the performance of Model-X Split Knockoffs when the number of samples and the sparsity level varies. In this figure, we adopt the bootstrap+ constructions of the Model-X Split Knockoff copies. Again, for all cases, the FDR of Model-X Split Knockoffs is under control. For each sparsity level, the selection power of Model-X Split Knockoffs gradually increases when the sample size increases.
@@ 5 Alzheimer’s Disease
5.1 Dataset Description
ADNI  [^6] http://adni.loni.ucla.edu dataset is collected by the structural Magnetic Resonance Imaging (MRI), which consists of 752 samples divided into three types: 126 AD, 433 Mild Cognitive Impairment(MCI), and 193 Normal Controls (NC). Each image is pre-processed by the toolbox Statistical Parametric Mapping (SPM) for segmentation into gray matter(GM), white matter (WM), and cerebral spinal fluid (CSF), followed by Automatic Anatomical Labeling(AAL) to partition the whole brain into 90 regions with the volume of each region.
To utilize the dataset, we construct a design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$, where $n=752$ and $p=90$, representing the volumes of $90$ processed brain regions for all $752$ samples. Specifically, $\mathbf{X}_{i,j}$ denotes the column-wise normalized volume of region $j$ for sample $i$. In this section, $\mathbf{Y}\in\mathbb{R}^{n}$ is taken as the Alzheimer’s Disease Assessment Scale (ADAS), a measure initially developed to evaluate the severity of cognitive dysfunction (Rosen et al., 1984), which was later shown to effectively distinguish between clinically diagnosed Alzheimer’s disease and normal controls (Zec et al., 1992).
@@ 5.2 Region Selection
Region Model-X Knockoff Model-X Split Knockoff
Frontal Inf Orb L
Frontal Inf Oper L
Parietal Inf R
Frontal Mid R
Rolandic Oper L
Rolandic Oper R
Cingulum Post L
Cingulum Post R
Hippocampus L
Hippocampus R
Lingual L
Precuneus R
Fusiform R
Supramarginal R
Temporal Mid L
Temporal Mid R
The region selection results of Model-X Split Knockoff and Model-X Knockoff are shown in Table 1. For Model-X Split Knockoff, the $W$ statistics is computed based on the cross-validated $\lambda$ and $\nu$ on a grid of $\log{\lambda}$ between $-2$ and $-1$ with a step size of $0.05$ and a grid of $\log{\lambda}$ between 00 and $1$ with a step size of $0.05$.
Overall, Model-X Split Knockoff identifies 15 brain regions associated with Alzheimer’s disease, whereas the standard Model-X Knockoff selects only 8, demonstrating the enhanced selection power of our approach.
Both methods select the hippocampus, middle temporal lobes (”Temporal Mid”), middle frontal gyrus (”Frontal Mid R”), and several cerebral cortex areas such as the lingual gyrus and fusiform gyrus. Most of these regions have been previously reported as potential biomarkers for Alzheimer’s disease in the literature (Lee et al., 2020, Peters et al., 2009, Yang et al., 2019). Specifically, the hippocampus and middle temporal lobes are known to be involved in memory function (Vemuri and Jack, 2010) and are well-established biomarkers in the progression of Alzheimer’s disease (Lee et al., 2020). Additionally, Yang et al. (2019) reported a significant reduction in cortical thickness in AD patients compared to normal controls (NC) across cerebral cortical regions, including the lingual gyrus and fusiform gyrus. Besides, Busatto et al. (2008) found that some sub-regions in the middle frontal gyrus suffered from degeneration in the early stage of the disease. These regions may exhibit strong discrimination in our dataset, causing our methods to favor this selection.
Besides, most of the regions that are additionally selected by our method have also been reported to undergo early degeneration in previous studies (Van Hoesen et al., 2000, Cajanus et al., 2019, Pengas et al., 2010, Lee et al., 2020). In particular, the inferior frontal gyrus plays multiple roles, including speech and language processing (Van Hoesen et al., 2000, Cajanus et al., 2019). Additionally, significant structural changes in nodal centrality within the rolandic operculum have been observed in the AD population (Yao et al., 2010, Bajo et al., 2015). Besides, the posterior cingulate cortex, which is involved in episodic memory, was found to exhibit atrophy at the earliest clinical stages of sporadic AD (Pengas et al., 2010, Lee et al., 2020).
@@ 5.3 Connection Selection
In this section, we detect abnormal connections between brain regions. We represent brain connectivity as a graph $G=(V,E)$, where $V$ denotes the brain regions and $E$ represents edges connecting adjacent regions. Based on this formulation, the transformational matrix $D$ is defined as a graph difference operator, such that $D(\beta)(i,j)=\beta_{i}-\beta_{j}$ for adjacent brain regions $(i,j)$. Since adjacent brain regions typically show similar activity, a significant signal on an edge suggests connections between relatively stable regions and severely atrophied regions contributing to the disease.
@@
Region 1 Region 2 Model-X Split Knockoff
Hippocampus L Cingulum Post L
Hippocampus L Fusiform L
Hippocampus L Precuneus L
Hippocampus L Putamen L
Hippocampus L Thalamus L
Hippocampus L Heschl L
Hippocampus R Lingual R
Hippocampus R Putamen R
Hippocampus R Temporal Sup R
Hippocampus R Temporal Inf R
Hippocampus R Insula R
Amygdala L Frontal Inf Orb L
Amygdala R Temporal Pole Sup R
Lingual L Fusiform L
Lingual L Thalamus L
Lingual L Cingulum Post L
Lingual R Occipital Mid R
Lingual R Fusiform R
Lingual R ParaHippocampal R
The abnormal connections identified by our method are presented in Table 2. Notably, many of the detected connections involve the hippocampus, which exhibits the earliest and most significant atrophy in Alzheimer’s disease (AD) (Greicius et al., 2003, Yang et al., 2019, Cajanus et al., 2019, Yao et al., 2010).
In addition, we identify two selected connections involving the amygdala. Anatomically adjacent to the hippocampus, the amygdala is also part of the medial temporal lobe and plays a key role in memory processing (Li et al., 2023). Notably, Nelson et al. (2018) reported that the amygdala is a primary site of amyloid plaque deposition in the early stages of AD—a hallmark that has been widely hypothesized as a defining feature of AD. Consistently, the amygdala has been shown to undergo early degeneration during disease progression (Ferry, 2012, Stouffer et al., 2023).
Moreover, the lingual regions were frequently involved in the selected connections. This finding is consistent with previous studies showing that the lingual gyrus—associated with visual processing and visual memory—plays an important role in the early detection of Alzheimer’s disease. Notably, impairments in lingual function have been suggested as another early indicator of AD (Yang et al., 2019, Van Hoesen et al., 2000).

### v2

@@ 4 Simulation Experiments
In this section, we evaluate the Model-X Split Knockoff method in two canonical settings, the normal distribution setting and the pairwise comparisons setting. The simulation results show that the Split Knockoff method achieves the desired false discovery rate control as well as relatively high selection power, specifically higher compared with Model-X Knockoff when both are applicable.
@@ 4.1 Normal Distribuions
Experimental Setting: In this section, simulation experiments are conducted under the case where the distribution of the design $X$ follows the normal distribution. In particular, rows of the design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ is generated independent and identically distributed (i.i.d.) from a multivariate normal distribution $\mathcal{N}(0_{p},{\bm{\Sigma}})$, where the covariance matrix $\bm{\Sigma}$} has the correlation structure ${\bm{\Sigma}}_{i,j}=c^{|i-j|}$ for all $(i,j)$ pairs with correlation parameter $c=0.5$. The true coefficient vector ${\beta}^{*}\in\mathbb{R}^{p}$ is defined by: $\beta^{*}_{i}=A$ if $i\leq k,i\equiv 0,-1\pmod{3}$; and $\beta^{*}_{i}=0$ otherwise. In this section, we consider the logistic regression, where the response $\mathbf{Y}\in\mathbb{R}^{n}$ is generated from $\mathbf{Y}\sim\text{Bernoulli}\left\{\sigma(\mathbf{X}\beta^{*})\right\}$, where $\sigma(x):=\frac{1}{1+\exp(-x)}$.
@@
The transformational sparsity is specified by the linear transformation matrix $D\in\mathbb{R}^{m\times p}$ such that $\gamma^{*}=D\beta^{*}$ is sparse. Based on the above specific choice of $\beta^{*}$, the following three kinds of $D$ are considered in this simulation:
  • $\beta^{*}$ itself is sparse, so that one can take $D_{1}=I_{p}$, leading to $m=p$.
@@
  • $\beta^{*}$ is a uni-dimensional piecewise constant function, so that one can take $D_{2}\in\mathbb{R}^{(p-1)\times p}$ as: $D_{2}(i,i)=1$, $D_{2}(i,i+1)=-1$ for $i\in\{1,\dots,p-1\}$, and $D_{2}(i,j)=0$ for other pairs of $(i,j)$. In this case, m=p1<pm=p-1<p.
@@
  • Combining the above two cases, one can take $D_{3}=\left(D^{\top}_{1},D^{\top}_{2}\right)^{\top}\in\mathbb{R}^{(2p-1)\times p}$, where m=2p1>pm=2p-1>p.
@@
Throughout this section, the target FDR level $q$ is set to be $0.2$, while the coefficients in the above simulation design are taken to be $p=100$, $k=20$, and $A=1$. For the case where $D_{1}=I_{p}$, the performance of Model-X Knockoffs [^5] The Model-X Knockoff method is implemented using the official repository. is also evaluated for comparisons. The Model-X (Split) Knockoff copies are generated from the data matrix $\mathbf{X}$ through the estimated covariance matrix $\widehat{\bm{\Sigma}}=\mathbf{X}^{\top}\mathbf{X}/n$. Specifically, the Split Knockoff copies $A$ and $\widetilde{A}$ are generated through Equation (16).
(i) $D=D_{1}$ (MK)
(ii) $D=D_{1}$ (MSK)
(iii) $D=D_{2}$ (MSK)
(iv) $D=D_{3}$ (MSK)
Figure 1: Performance of Model-X Knockoffs (when applicable) and Model-X Split Knockoffs when $q=0.2$ and the number of samples varies. The curves indicate the mean performance, with the shaded areas representing the standard deviation across 200 simulations. For brevity, we denote Model-X Knockoffs and Model-X Split Knockoffs as MK and MSK, respectively.
@@
In the first two figures of Figure 1, we compare the performance of Model-X Knockoffs and Model-X Split Knockoffs when $D=D_{1}=I_{p}$. In both figures, with the increase of sample size, the selection power of both methods gradually increases from 0 to 1, while the FDR of both methods are under control for all sample sizes. In terms of selection power, the Model-X Split Knockoff demonstrably yields superior results across all tested sample sizes, a finding supported by Proposition 2.1. We hypothesize that this performance gain is a direct consequence of the improved incoherence conditions, which arise from the orthogonal design of $A$ and $\widetilde{A}$ in Equations (16).
@@
In the last two figures of Figure 1, the performance of Model-X Split Knockoffs is evaluated in the scenarios where $D$ is non-trivial, i.e., $D\neq I_{p}$. In both cases, the FDR of Model-X Split Knockoffs (MSK+) is under control universally for all choices of sample sizes, with MSK higher as defined. Meanwhile, the selection power of Model-X Split Knockoffs follows a similar trend as the case where $D=I_{p}$, i.e., the selection power gradually increases from 0 to 1 when the sample size enlarges.
@@ 4.2 Pairwise Comparisons
Experimental Settings: In this section, the pairwise comparison problem among $p$ objects with $n$ noisy observations of comparisons is considered. The scores for $p$ objects, denoted as $\beta^{*}\in\mathbb{R}^{p}$, are generated from a standard normal distribution $\mathcal{N}(0,1)$. To introduce sparsity into the dataset, the scores in $\beta^{*}$ are truncated according to a sparsity ratio $k$ set at 0.5. Specifically, only the first $k\cdot p$ components of $\beta^{*}$ are retained, while the remaining components are set to zero.
The rows of the pairwise comparison design matrix $\mathbf{X}\in\{0,1,-1\}^{n\times p}$ are generated independently. For the $i$-th row, two indices $1\leq j<k\leq p$ are chosen uniformly at random with probability $\frac{2}{p(p-1)}$ for each possible pair of $(j,k)$. Then we set $\mathbf{X}_{i,j}=1$, $\mathbf{X}_{i,k}=-1$, and $\mathbf{X}_{i,l}=0$ for $l\neq j,k$. This specific construction ensures that each row of $\mathbf{X}$ represents a (noisy) comparison between two objects.
In this simulation experiment, the performance of Model-X Split Knockoff is evaluated under the Bradley-Terry model, where the response vector $\mathbf{Y}\in\mathbb{R}^{n}$ is generated by the logistic regression $\mathbf{Y}\sim\text{Bernoulli}\left\{\sigma(\mathbf{X}\beta^{*})\right\}$. The transformational matrix $D\in\mathbb{R}^{\frac{p(p-1)}{2}\times p}$ is taken to be the graph difference operator on a fully connected graph with $p$ vertices to conduct pairwise comparisons.
@@
(iii) Runtime (Second)
Figure 2: Performance and runtime of Model-X Split Knockoff when $q=0.2$ as the number of features varies for bootstrap+ and sequential constructions. The curves indicate the mean performance, with the shaded areas representing the standard deviation across 200 simulations.
Figure 2 compares the performance of bootstrap+ constructions of Model-X Split Knockoff copies against the sequential constructions. It shows that, although the two types of constructions achieve desired FDR control, the selection power of the bootstrap+ constructions is higher compared with the sequential constructions for all choices of the number of features. This is explained in Section 2.2.2 that the bootstrap+ constructions makes sure $\widetilde{X}^{A}$ to be independent from $A$, which helps improve the incoherence conditions and thus the selection power.
Moreover, when $p$ is large, the sequential construction may need to estimate too many conditional distributions, so the computation cost explodes. On the other hand, the bootstrap+ constructions do not suffer from this issue and handle such a case easily. As shown in Figure 2, the sequential method exhibits a much steeper increase in running time compared to the bootstrap+ method, which requires significantly less computation time under the same settings.
Due to the above two points, the bootstrap+ construction will be the default construction to implement Model-X Split Knockoff for pairwise comparisons.
@@ 5 Alzheimer’s Disease
We apply our method to the structural Magnetic Resonance Imaging (MRI) in the ADNI dataset. We extracted 752 samples that are composed of 126 AD, 433 Mild Cognitive Impairment (MCI), and 193 Normal Controls (NC). Applying Automatic Anatomical Labeling (AAL), we extract the volumes of $90$ processed brain regions for all $752$ samples. Therefore, for our design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$, $\mathbf{X}_{i,j}$ denotes the column-wise normalized volume of region $j$ for sample $i$. Besides, $\mathbf{Y}\in\mathbb{R}^{n}$ is taken as the Alzheimer’s Disease Assessment Scale (ADAS), a measure initially developed to evaluate the severity of cognitive dysfunction (Rosen et al., 1984), which was later shown to effectively distinguish between clinically diagnosed Alzheimer’s disease and normal controls (Zec et al., 1992).
@@ 5.1 Region Selection
The region selection results of Model-X Split Knockoff and Model-X Knockoff are shown in Table 1. Overall, Model-X Split Knockoff identifies 12 brain regions associated with Alzheimer’s disease, whereas the standard Model-X Knockoff selects only 8, demonstrating the enhanced selection power of our approach.
@@
Region IPL R ORBsup L MFG R ROL L SFGmed R PCG L PCG R HIP L HIP R LING L FFG R MTG L MTG R ITG L
Model-X Knockoff
Model-X Split Knockoff
Both methods select the hippocampus (“HIP L”, “HIP R”), middle temporal lobes (“MTG L”, “MTG R”), middle frontal gyrus (“MFG R”), and several cerebral cortex areas such as the lingual gyrus (“LING L”, “LING G”) and fusiform gyrus (“FFG R”). These regions have been previously reported as potential biomarkers for Alzheimer’s disease in the literature (Lee et al., 2020, Peters et al., 2009, Yang et al., 2019). Most of additional regions identified by our method have been reported to show early degeneration in previous studies (Van Hoesen et al., 2000, Cajanus et al., 2019, Pengas et al., 2010, Lee et al., 2020), with the exception of the Superior frontal gyrus, orbital part (“ORBsup L”), which may represent a false discovery. In particular, significant structural changes in nodal centrality within the rolandic operculum (“ROL L”) have been observed in the AD population (Yao et al., 2010, Bajo et al., 2015). Besides, the posterior cingulate cortex (“PCG L”, “PCG R”), which is involved in episodic memory, was found to exhibit atrophy at the earliest clinical stages of sporadic AD (Lee et al., 2020).
@@ 5.2 Connection Selection
In this section, we detect abnormal connections between brain regions. We represent brain connectivity as a graph $G=(V,E)$, where $V$ denotes the brain regions and $E$ represents edges connecting adjacent regions. Based on this formulation, the transformational matrix $D$ is defined as a graph difference operator, such that $(D\beta)_{(i,j)}=\beta_{i}-\beta_{j}$ for adjacent brain regions $(i,j)$. Since adjacent brain regions typically show similar activity, a significant signal on an edge suggests connections between relatively stable regions and severely atrophied regions contributing to the disease. The abnormal connections identified by our method are presented in Table 2.
@@
HIP L & PHG L HIP L & LING L HIP L & FFG L HIP L & PUT L
HIP L & HES L HIP R & PHG R HIP R & TPOmid R HIP R & ITG R
HIP L & INS L LING L & FFG L LING R & PHG R ROL L & HES L
MTG L & INS L MTG L & MOG L MTG R & TPOsup R PCG R & PCUN R
PCL R & IPL R
Notably, many of the detected connections involve the hippocampus (“HIP L”, “HIP R”), which exhibits the earliest and most significant atrophy in Alzheimer’s disease (AD) (Greicius et al., 2003, Yang et al., 2019, Cajanus et al., 2019, Yao et al., 2010).
In addition, some connections involve the middle temporal lobes (“MTG L”, “MTG R”). The middle temporal lobe, a key component of the temporal cortex, plays an essential role in semantic memory and language processing, and is implicated in the early pathological changes of Alzheimer’s disease (AD) (Chen et al., 2022). Moreover, the lingual regions (“LING L”, “LING R”) were involved in the selected connections. This finding is consistent with previous studies showing that the lingual gyrus—associated with visual processing and visual memory—is important in the early detection of Alzheimer’s disease.